# Robust Utility-Preserving Text Anonymization Based on Large Language Models

### v1

@@ 3.4 Lexicographic Optimizer
This iterative process continues until either the pre-defined maximum values for both objectives are reached or the maximum number of iterations $T$ is met. By continuously refining and evaluating the anonymized text, the optimizer iteratively improves it to achieve an optimal balance between privacy and utility.
3.5 Distilling the Anonymization Ability
Utilizing LLMs for text anonymization is computationally expensive, and for certain LLMs, access is only available through APIs, which raises privacy and cost concerns. However, in our framework, the optimization result heavily depends on the reasoning ability of the LLM, which stems from the large scale of parameters these models possess. Recent studies have demonstrated that prompting LLMs as optimizers is less effective with smaller-scale models Zhang et al. (2024).
To address this issue, we employ knowledge distillation (KD), where a large model (the teacher) transfers its knowledge to a smaller model (the student). Typically, KD involves training the student model using the outputs of the teacher model as labels Kim and Rush (2016). In our case, we utilize the final anonymization result produced by the teacher model during the lexicographic optimization as the training label for the student model.
To utilize the generation results of the teacher model more efficiently, we adopt the Direct Preference Optimization (DPO) Rafailov et al. (2023) method. This method fine-tunes an LLM on human labels of the relative quality of model generations to align the model with human preferences. In our method, intermediate optimization results from the teacher model can be considered less preferred than the final optimization result. These intermediate and final results form the preference dataset. We fine-tune the student model using DPO on this dataset to preferentially generate outputs similar to the final optimization result while reducing the likelihood of producing results akin to the intermediate stages.
@@ 4 Experimental Set-up
Dataset #Train #Validation #Test
DBPedia Classes 1938 243 239
Personal Reddit 318 - 207
Table 1: Statistics of experiment datasets.
Method Disclosure Risk Utility Preserving
@@
RUPTA (Llama-3-70b) 64.02 63.23 95.34 96.23 95.55 95.82 0.2224
RUPTA (GPT-3.5) 68.51 69.16 95.40 96.02 95.70 95.49 0.2188
RUPTA (GPT-4) 52.67 53.11 95.58 96.26 95.91 96.02 0.1618
@@
Table 2: Main experiment results on the test set of DB-bio dataset. The top and second performance are highlighted with bold font and underline, respectively.
Datasets.
We evaluate our text anonymization method on the following two datasets:
  • We sampled celebrity biographies from the DBpedia Classes dataset Dan (2019) to build a new dataset DB-bio. Unlike the commonly-used Wiki-bio dataset Lebret et al. (2016) in anonymization studies that lacks annotations for downstream tasks, this dataset includes detailed three-level hierarchical category annotations. We use the third-level category labels as occupation classification labels to assess the impact of our anonymization method on this specific downstream task. The name of the person described by the biography is used as the ground-truth personal information.
  • To further validate the generality of our method, we evaluate it on the PersonalReddit (PR) dataset  Staab et al. (2024a) consisting of 525 human-verified synthetic public Reddit comments and corresponding user profiles. We use the annotated occupation attribute in the profile as the label of the occupation classification task and anonymize the comments to prevent the identification of other personal attributes.
General statistics of these datasets can be seen in table 1. Detailed statistics, including category distributions, are provided in appendix A.
Evaluation Metrics.
To evaluate our text anonymization method, we focus on two critical aspects: disclosure risk and utility preservation. Disclosure risk is assessed by measuring the Success Rate (SR) of a strong adversarial LLM in inferring personal information from anonymized text. Additionally, we prompted an LLM to generate the Confidence Scores (CS), evaluating the degree of confidence with which anonymized text can be linked to the ground-truth personal information.
Utility preservation metrics are gauged by the performance of a simple neural network classifier trained on non-anonymized train data and tested on anonymized text, including Accuracy, Precision, Recall, F1 Score, and the classifier’s loss function value indicating classification uncertainty. Specific metric settings can be seen in appendix B.
Comparison Methods.
To establish the effectiveness of our text anonymization framework, we benchmark it against state-of-the-art methods and industry standards.
  • We use Azure Aahill (2023)’s industry-standard state-of-the-art text anonymizer as a traditional anonymization baseline.
  • AF Staab et al. (2024b) is a current state-of-the-art method for text anonymization based on the adversarial feedback mechanism.
  • DEID-GPT Liu et al. (2023) prompts the LLM to mask out all the entities of pre-defined kinds.
  • SD Dou et al. (2023) prompts the LLM to replace the entities of pre-defined kinds with more general counterparts.
All these methods are recreated using the GPT-4 model Achiam et al. (2023). Besides, we explore the effectiveness of using different LLM architectures as the lexicographic optimizer, including open-sourced models like instruction-tuned Llama-3-70b AI@Meta (2024) and Mixtral $8\times 22$Jiang et al. (2024), and the proprietary GPT-4 and GPT-3.5. Besides, we evaluate the original non-anonymized dataset (Original) for reference.
Implementation Details.
GPT-4 is used exclusively as the privacy evaluator of RUPTA and simulated attacker of AF due to its advanced capabilities in re-identification. GPT-4 is also used as the utility evaluator of RUPTA. Besides, we experimented with using Phi-3 Mini Abdin et al. (2024) and Llama-3-8b AI@Meta (2024) as the student model. Details can be seen in appendix C.
5 Experimental Results
5.1 Overall Results
Method Disclosure Risk Utility Preserving
[…]
Personal Reddit Original 49.76 81.89 55.13 63.51 55.80 58.45 1.5695
Azure Aahill (2023) 45.89 81.07 54.04 58.49 54.17 57.00 1.7340
DEID-GPT Liu et al. (2023) 43.12 72.81 53.98 58.21 54.06 56.31 1.9314
SD Dou et al. (2023) 44.05 75.17 54.11 58.43 54.21 56.93 1.7501
AF Staab et al. (2024b) 35.40 57.76 16.64 22.32 16.68 21.26 3.3380
RUPTA (Llama-3-70b) 39.61 61.63 32.96 44.57 32.82 38.65 2.3131
RUPTA (GPT-3.5) 34.30 61.50 32.04 40.44 31.97 36.23 2.4477
RUPTA (GPT-4) 35.75 55.04 30.34 39.14 30.09 35.75 2.5391
Table 3: Experimental results on the test set of PersonalReddit dataset. The top and second performance are highlighted with bold font and underline, respectively.
The overall experimental results on the DB-bio dataset are presented in table 2. In the disclosure risk evaluation, methods that anonymize the data in an iterative refinement manner, including our RUPTA method and the AF method, achieve the best performance. Although DEID-GPT and SD also leverage LLMs, they follow a traditional approach focusing on masking entities of pre-defined types. Experiment results demonstrate that such methods cannot adequately defend against re-identification attacks from LLMs. Additionally, using open-source LLMs as the lexicographic optimizer also achieves comparable privacy-preserving performance, demonstrating the practicality and generality of our method.
For the utility preserving evaluation, traditional methods like Azure mask all the entities of pre-defined kinds with “*”, leading to the most significant information loss, thus achieving the lowest performance. The DEID-GPT, SD, and AF methods, although anonymized through replacing sensitive entities with more general ones, do not consider the downstream analysis task and generalize all the possible sensitive entities, which also significantly undermines downstream task performance. Visualization results of the optimization process in fig. 3 highlight the drawback of the AF method, where the SR and classification accuracy decrease simultaneously as the number of optimization steps increases. In contrast, our method achieves the best downstream task performance. Furthmore, during the optimization process of RUPTA, there is an explicit increasing phase of the classification accuracy, demonstrating the effectiveness of the RUPTA method to maximize both the privacy and utility in the anonymization process. This trend also illustrates that beyond a certain point, further anonymization yields diminishing returns in privacy preservation and results in greater losses of utility information.
5.2 Customizable Privacy-Utility Tradeoff
Figure 4: Customizable privacy-utility tradeoff experiments on the test set of DB-bio dataset with GPT-4, Llama-3-70b (Llama-3), and Mixtral $8\times 22$b (Mixtral) as optimizers, respectively.
The experiment results for the customizable privacy-utility tradeoff are displayed in fig. 4. In our method, the maximum value of the privacy objective is set manually according to specific requirements, allowing for a customizable privacy-utility tradeoff. We analyze and visualize the average SR and classification accuracy of our method using GPT-4, Llama-3-70b, and Mixtral $8\times 22$b as the lexicographic optimizer. We set the maximum privacy value to 1, 5, 10, 15, and 20, respectively. It is evident in fig. 4 that our proposed method can effectively adapt the privacy-preserving level according to the maximum value setting. As the maximum privacy value increases, the average privacy score improves while the utility score adjusts accordingly. This observation demonstrates the flexibility of our approach in balancing privacy and utility based on user-defined requirements.
5.3 Experiments on PR Dataset
To evaluate the generality of our method, we further conduct experiments on the PR dataset with results presented in table 3. The PR dataset is characterized by fewer explicit and more implicit sensitive entities. Entity recognition-based methods, including Azure, DEID-GPT, and SD, struggle to detect these implicit entities, resulting in minimal masking operations, as evidenced by their evaluation results closely mirroring those of the original dataset. Consequently, while these methods exhibit higher performance on the downstream task, they provide inferior privacy protection. Only the AF and our method can properly detect implicit sensitive information and achieve the lowest disclosure risk. However, the AF method anonymizes without tailoring its approach to the specific downstream task, which significantly impairs task performance. In contrast, our method not only effectively minimizes disclosure risk but also preserves a greater degree of utility in the anonymized text than AF, achieving a better privacy-utility tradeoff.
5.4 Distilled Models
In this experiment, we try to distill the anonymization ability of GPT-4 into lightweight models. Using RUPTA with GPT-4 as the lexicographic optimizer, we anonymized the training and validation sets of the DB-bio dataset. Initially, we fine-tuned student models in a supervised manner (SFT) using the final optimization results as labels. Then we constructed a preference dataset from the optimization trajectories and conducted DPO fine-tuning on the optimal checkpoint during the SFT phase. The evaluation results are presented in fig. 5.
Figure 5: Knowledge distillation experiment results using Llama-3-8b (Llama-3) and Phi-3 Mini (Phi-3) as the student model, respectively.
In the utility preserving evaluation results, in addition to the classification accuracy, we further demonstrate the semantic similarity between the anonymized and original text. The supervised fine-tuned student models maintain a high level of downstream task performance. Although the DPO fine-tuning process improves the privacy-preserving performance, it somewhat harms the downstream task performance. This phenomenon likely results from the unbalanced optimization phases in the lexicographic optimization process, where achieving the maximum privacy objective requires more iterations than improving downstream task performance, as shown in fig. 3. Consequently, the student models, fine-tuned with DPO, prioritize privacy to a greater extent, potentially at the expense of utility. Anonymization examples are shown in fig. 6. We can see that the student model can learn to generalize or remove sensitive entities after the SFT phase. After the DPO fine-tuning phase, the student model can further generalize sensitive entities marked by underlining, e.g., from “father” to “family member”. Both models can keep the relevant information about the downstream task in the anonymized text, as highlighted in the figure.
Figure 6: Anonymization example of Phi-3 Mini model
6 Conclusions
This paper presents a novel framework that integrates a privacy evaluator, a utility evaluator, and an optimizer to effectively anonymize text for text anonymization using LLMs, ensuring reduced risk of re-identification while maintaining utility for downstream tasks. Building on that, we further develop practical methods based on DPO to distill the anonymization capabilities into lightweight models with the performance comparable to that of the teacher models. Additionally, the creation of a new dataset based on celebrity biographies with occupation labels provides a valuable resource for assessing the impact of various anonymization techniques on the specific downstream task-occupation classification. The superiority of our methods over existing models contributes to text anonymization and sets new baselines for future research that considers downstream utility in anonymization.
Limitations
While our study presents significant advancements in text anonymization techniques using LLMs, there are several limitations to acknowledge and to be mitigated in the future work.
Ethics Statement
This research adheres to ethical guidelines in the development and application of text anonymization technologies using LLMs. Recognizing the dual-edged nature of anonymization—its potential to protect privacy while also possibly enabling data misuse—we have implemented several safeguards to ensure responsible use. We commit to transparency in our methodologies and the limitations of our models, as detailed in previous sections of this paper. By openly discussing the strengths and weaknesses of our approach, we aim to foster an informed community that can critically assess and improve upon our work. Besides, while developing our dataset from celebrity biographies, we have ensured that all data used were sourced from publicly available, non-sensitive information. The dataset complies with all applicable data protection laws and ethical standards, and no personally identifiable information was used without consent.
Acknowledgement
Appendix A Dataset Settings

### v2

@@ 3.4 The Optimizer
The iterative process continues until either the pre-defined maximum values for both objectives are reached or the maximum number of iterations $T$ is met.
3.5 Distilling Anonymization Ability
Iterative anonymization methods based on LLMs could be time-consuming and resource-intensive. We investigate the sequence-level knowledge distillation (SKD) Kim and Rush (2016), where a large model (the teacher) transfers its knowledge to a smaller model (the student). Specifically, we propose to utilize the final anonymization result produced by the teacher model during the lexicographic optimization as the training label for the student model.
To utilize the generation results of the teacher model more efficiently, we adopt the Direct Preference Optimization (DPO) Rafailov et al. (2023) method. This method fine-tunes an LLM on human labels of the relative quality of model generations to align the model with human preferences. In our method, intermediate optimization results from the teacher model can be assumed less preferred than the final optimization result. These intermediate and final results form the preference dataset. We fine-tune the student model using the DPO method on this dataset to preferentially generate outputs similar to the final optimization result while reducing the likelihood of producing results akin to the intermediate stages.
@@ 4 Experimental Set-up
Method Disclosure Risk Utility Preservation
[…]
DB-bio Original 100.00 98.45 99.58 99.68 99.61 99.58 0.0422
Azure Aahill (2023) 78.24 80.87 91.63 95.04 92.39 92.47 0.3202
DEID-GPT Liu et al. (2023) 77.10 79.47 90.82 94.37 92.56 91.22 0.3103
SD Dou et al. (2023) 73.21 73.63 92.27 93.11 92.69 92.96 0.2719
IncogniText Frikha et al. (2024) 58.06 56.28 85.68 89.03 87.32 88.28 0.4842
AF Staab et al. (2024b) 52.91 50.84 91.20 94.26 91.75 92.02 0.4048
RUPTA (Llama-3-70b) 64.02 63.23 95.34 96.23 95.55 95.82 0.2224
RUPTA (GPT-3.5) 68.51 69.16 95.40 96.02 95.70 95.49 0.2188
[…]
Table 1: The main experiment results on the test set of the DB-bio dataset. The top and second performances are highlighted with bold font and underlined, respectively. Results of RUPTA (GPT-4) denoted by $\dagger$ are significantly better than that of the AF method under the one-tailed paired t-test $(p<0.05)$.
Datasets.
Following previous studies Staab et al. (2024a); Morris et al. (2022), we evaluate our model on the DB-bio and PersonalReddit datasets. We investigate the impact of anonymization methods on the occupation classification task, a real-world task involving personal information De-Arteaga et al. (2019). Occupation classification is critical for applications such as job recommendation platforms and automated hiring tools. In such contexts, anonymizing other personal information without affecting the occupation classification performance is important. In addition, a simple classification task allows us to easily demonstrate the performance of the models through quantitative and qualitative analysis. It helps provide intuitive insights into how anonymization models alter critical words necessary for accurate classification.
  • DB-bio: Previous anonymization studies Morris et al. (2022) have been conducted on celebrity data available in Wikipedia. Inspired by that, we sampled celebrity biographies from the DBpedia Classes dataset Dan (2019) to build a new DB-bio dataset for our study and future research, where we used the category labels of each celebrity in the DBpedia Classes as the occupation classification label.
  • PersonalReddit (PR) Staab et al. (2024a): To further assess our method, we evaluate it on the PersonalReddit dataset where we assume personal attributes like gender and location as sensitive information and we use occupations as the labels of the task.
Detailed statistics, including category distributions, are provided in appendix D.
Evaluation Metrics.
The evaluation focuses on two critical aspects: disclosure risk and utility preservation. Disclosure risk is assessed by measuring the Success Rate (SR) of a state-of-the-art LLM in inferring personal information from anonymized text. Besides, we prompted the LLM to generate the Confidence Scores (CS), evaluating the degree of confidence with which anonymized text can be linked to the ground-truth personal information.
Utility preservation metrics are gauged by the performance of a BERT-based classifier finetuned on original train data and tested on the test data that is anonymized by RUPTA and other methods, including Accuracy, Precision, Recall, F1 Score, and the classifier’s loss function value indicating classification uncertainty. More details and discussion about metrics can be seen in appendix F.
Models in Comparison.
We compare RUPTA with the following state-of-the-art models.
  • We include an industry-standard text anonymizer from Microsoft Azure Aahill (2023) as one of our anonymization baselines.
  • AF Staab et al. (2024b) is a current state-of-the-art LLM-based method for text anonymization based on an adversarial feedback mechanism.
  • DEID-GPT Liu et al. (2023) is a recent model that prompts LLMs to mask out pre-defined types of entities.
  • SD Dou et al. (2023) is another state-of-the-art approach prompting LLMs to replace entities with more general concepts.
  • IncogniText Frikha et al. (2024) is another LLM-based anonymization method that iteratively rewrites the text according to adversarial feedbacks. Different from AF, this method adds synthetic information during rewriting to mislead the attacker and proposes an early stopping mechanism to preserve utility.
The LLMs used in the above models are the state-of-the-art GPT-4 models Achiam et al. (2023). In addition, we explore the effectiveness of different LLMs for the lexicographic optimizer, including open-sourced Llama-3-70b AI@Meta (2024) and Mixtral $8\times 22$Jiang et al. (2024), and the proprietary GPT-4 and GPT-3.5 OpenAI (2022).
Implementation Details.
We use GPT-4 as our backbone LLM to ensure the model is comparable to the baselines. The original non-anonymized dataset is evaluated (Original) for reference. For implementation details including those for the distillation models, please refer to appendix G.
5 Experimental Results
5.1 Overall Performance
Method Disclosure Risk Utility Preservation
[…]
Personal Reddit Original 49.76 81.89 55.13 63.51 55.80 58.45 1.5695
Azure Aahill (2023) 45.89 81.07 54.04 58.49 54.17 57.00 1.7340
DEID-GPT Liu et al. (2023) 43.12 72.81 53.98 58.21 54.06 56.31 1.9314
SD Dou et al. (2023) 44.05 75.17 54.11 58.43 54.21 56.93 1.7501
IncogniText Frikha et al. (2024) 37.55 60.02 10.19 11.06 10.61 13.47 4.4766
AF Staab et al. (2024b) 35.40 57.76 16.64 22.32 16.68 21.26 3.3380
RUPTA (Llama-3-70b) 39.61 61.63 32.96 44.57 32.82 38.65 2.3131
RUPTA (GPT-3.5) 34.30 61.50 32.04 40.44 31.97 36.23 2.4477
RUPTA (GPT-4) 35.75 55.04 $30.34^{\dagger}$ $39.14^{\dagger}$ $30.09^{\dagger}$ $35.75^{\dagger}$ $2.5391^{\dagger}$
Table 2: Experimental results on the test set of the PersonalReddit dataset. The top and second performances are highlighted with bold font and underlined, respectively. Results of RUPTA (GPT-4) denoted by $\dagger$ are significantly better than that of the AF method under the one-tailed paired t-test $(p<0.05)$.
The overall experimental results on the DB-bio dataset are presented in table 1. We can see that in the disclosure risk evaluation, methods that anonymize the data in an iterative refinement manner, including RUPTA, IncogniText and AF, achieve the best performance. Although DEID-GPT and SD also leverage LLMs, they follow a traditional approach focusing on masking entities of pre-defined types. Experiment results demonstrate that such methods are not able to adequately defend against re-identification attacks from LLMs, posing critical concerns to these privacy protection methods. Additionally, we can see that using open-source LLMs to build the optimizer can achieve privacy-preserving performance comparable to closed-source models, demonstrating the generality of our method.
RUPTA achieves the lowest utility loss in utility preservation evaluations compared to other baselines. Anonymization generally reduces data specificity to protect privacy, but this comes at the cost of reduced utility. While the original data provides high utility with minimal privacy protection, full masking maximizes privacy but significantly diminishes utility. It is thus essential to investigate a balance—reducing specificity to safeguard private information while preserving sufficient utility in downstream tasks. Existing baselines, however, only assess utility after anonymization and often fail to maintain this balance, either offering insufficient privacy protection (SD and DEID-GPT) or applying excessive anonymization that compromises utility (e.g., AF and IncogniText). Due to the introduction of distracting synthetic information, IncognitText gets the worst utility-preserving performance. RUPTA respects and achieves super disclosure-risk performances and maintains competitive utility preservation.
We hope this work helps set a new benchmark for text anonymization research in the context of LLMs and under the practical setup of examining both the protection and utility, since without an explicit consideration of the two perspectives, we lose a comprehensive view of the problem.
In fig. 3, the visualization of the evaluation results of the optimization process shows that the SR and classification accuracy decrease simultaneously as the number of optimization steps increases. In contrast, our method achieves the best performance in the downstream task. During the optimization process of RUPTA, there is an explicit increase phase of the classification accuracy, demonstrating the effectiveness of the RUPTA method to maximize both privacy and utility in the anonymization process. Similar trends can also be seen from the average $u_{t}$ score during the anonymization process on the PersonalReddit dataset, as shown in table 3. This trend illustrates that, beyond a certain point, further anonymization yields diminishing returns in privacy preservation and results in greater losses of utility information. Baseline methods that either ignore the downstream task utility during the anonymization process or only attempt to mitigate its loss afterward often miss this certain point. As a result, they may cause greater utility loss or improve utility at the expense of privacy protection performance. To further demonstrate the effectiveness of RUPTA, we conducted experiments that adapted DEID-GPT and SR as lexicographic optimizers, as detailed in appendix B.
$t$ 1 2 3 4
$u_{t}$ 43.14 42.31 43.30 44.08
Table 3: Average $u_{t}$ score during the anonymization process on the PersonalReddit dataset.
5.2 Customizable Privacy-Utility Tradeoff
The experiment results for the customizable privacy-utility tradeoff are displayed in fig. 4. In our method, the maximum value of the privacy objective $K+1$ is set manually according to specific requirements, allowing for a customizable privacy-utility tradeoff. We analyze and visualize the average SR and classification accuracy of our method using GPT-4, Llama-3-70b, and Mixtral $8\times 22$b as the lexicographic optimizer. We set the maximum privacy value to 1, 5, 10, 15, and 20, respectively.
Figure 4: Customizable privacy-utility tradeoff experiments on the test set of the DB-bio dataset with GPT-4, Llama-3-70b (Llama-3), and Mixtral $8\times 22$b (Mixtral) as optimizers, respectively.
It is evident in fig. 4 that our proposed method can effectively adapt the privacy protection level according to the maximum value setting. As the maximum privacy value increases, the average privacy score improves while the utility score adjusts accordingly. According to fig. 3, the privacy protection level also varies throughout the optimization process. Thus RUPTA can adjust the privacy protection level in a wider range by adjusting both the maximum number of iterations $T$ and $K$. More experiment results are included in appendix C.
5.3 Experiments on the PR Dataset
To demonstrate the generality of our method, we further conduct experiments on the PR dataset with results presented in table 2. The PR dataset is characterized by fewer explicit and more implicit sensitive entities. Entity recognition-based methods, including Azure, DEID-GPT, and SD, struggle to detect these implicit entities, resulting in minimal masking operations, as evidenced by their evaluation results closely mirroring those of the original dataset. Consequently, while these methods exhibit higher performance on the downstream task, they provide inferior privacy protection. Only the AF and RUPTA can properly detect implicit sensitive information and achieve the lowest disclosure risk. However, the AF method anonymizes without tailoring its approach to the specific downstream task, which significantly impairs task performance. In contrast, RUPTA not only effectively minimizes disclosure risk but also preserves a greater degree of utility in anonymized text than AF.
5.4 Distillation Results
As shown in the computational analysis in appendix A, anonymization methods based on iterative prompting of LLMs are computationally expensive. We follow the knowledge distillation scheme proposed in section 3.5 to distill the anonymization ability of GPT-4 into lightweight models. The evaluation results are presented in fig. 5. More detailed results can be seen in appendix E.
Figure 5: Results of the knowledge distillation experiment using Llama-3-8b (Llama-3) and Phi-3 Mini (Phi-3) as the student model, respectively.
In the utility preservation evaluation results, in addition to the classification accuracy, we further demonstrate the semantic similarity between the anonymized and original text. The supervised fine-tuned student models maintain a high level of downstream task performance. We observed that although the DPO fine-tuning process improves the privacy-preserving performance, it could harm the downstream task performance. We due this to the objective of the optimizer in prioritizing privacy over utility, where more optimization steps are utilized to improve the level of privacy protection as shown in fig. 3. The student models have been shown to learn from the optimization history data and prioritize privacy to a greater extent potentially at the expense of utility. Anonymization examples are shown in fig. 6. We can see that the student model can learn to generalize or remove sensitive entities after the SFT phase. After the DPO fine-tuning phase, the student model can further generalize sensitive entities marked by underlining, e.g., from “father” to “family member”. Both models can keep the relevant information in the anonymized text for the downstream task, as highlighted in the figure.
Figure 6: Anonymization example of Phi-3 Mini model
5.5 Human Evaluation
To further examine the semantic loss and the quality of generated text, we conducted a human evaluation on a subset of 100 randomly selected examples from the test split of DB-bio, where 3 non-author human subjects rated the anonymization output of Azure, SD, AF and RUPTA using a Likert scale of 1 - 5, on "how much the original meaning of the non-anonymized text is preserved after anonymization?" (5 means all information is preserved and 1 means none). In addition, we also asked the human subjects to evaluate the fluency of the anonymized text with the Likert scale of 1 - 5 (5 means the most fluent and 1 means the least). We take an average of the scores of the three human subjects. The results can be seen in table 4.
Method Semantic Preservation Fluency
Azure 3.23 2.09
SD 3.64 3.62
AF 3.78 3.71
RUPTA 3.96 3.68
Table 4: Human evaluation results on the test set of DB-bio.
We can see that Azure introduces the most significant distortion to the original text, as it masks sensitive entities with extensive nonsense special tokens. SD replaces sensitive entities with more general English terms, resulting in better overall semantic preservation and fluency performance. Unlike Azure and SD, which are based on entity-level replacement and adaptation, RUPTA and AF are rewriting-based methods that rewrite the entire passage, which makes them achieve similar or better fluency performance. Compared to the other three methods, RUPTA achieves the best semantic preservation performance, which we believe is due to its consideration of the general semantics of the entire passage during anonymization.
6 Conclusions
This paper presents a novel framework, RUPTA, that integrates a privacy evaluator, a utility evaluator, and an optimizer to effectively anonymize text, ensuring reduced risk of re-identification while maintaining utility for downstream tasks. Building on that, we further develop practical methods based on DPO to distill the anonymization capabilities into lightweight models with a performance comparable to that of the teacher models. Additionally, we create a dataset from celebrity biographies with occupation labels to evaluate the effects of anonymization techniques on specific tasks. The superior performance of RUPTA over existing models and the evaluation setup help establish new baselines for future research that considers downstream task utility in anonymization.
Limitations
While our study presents significant advancements in text anonymization techniques using LLMs, there are several limitations to acknowledge and to be mitigated in future work.
Ethics Statement
Recognizing the dual-edged nature of anonymization—its potential to protect privacy while also possibly enabling data misuse—we have implemented several safeguards to ensure responsible use. We commit to transparency in our methodologies and the limitations of our models, as detailed in previous sections of this paper. By openly discussing the strengths and weaknesses of our approach, we aim to foster an informed community that can critically assess and improve upon our work. Besides, this work is evaluated on publicly available datasets. While developing our dataset from celebrity biographies, we have ensured that all data used were sourced from publicly available, non-sensitive information.
Acknowledgement
Appendix A Computational Cost Analysis
To analyze the computational cost of anonymization method based on iterative prompting LLMs, including AF and RUPTA, we record the average time for anonymizing one paragraph in the DB-bio dataset (AT), average number of prompt tokens (PT) and average number of completion tokens (CT) in table 5. Each paragraph contains 234 tokens on average in this dataset.
Method AT (s) #PT #CT
RUPTA 76.38 3846.28 697.21
AF 72.57 2979.34 612.01
Table 5: Computational cost analysis of the LLM-based methods.
As shown by the results, the computational cost in time and money are both relatively high, which demonstrates the necessity of distilling the knowledge of a large model into a lightweight model to improve the practicality of these methods.
Appendix B Method Transferring
To further demonstrate the effectiveness and practicality of RUPTA, we conducted experiments about adapting previous LLM-based anonymization methods that anonymize by prompting the LLM in a single round including DEID-GPT and SD as the optimizer in RUPTA. Specifically, we append an additional prompt to make them conduct the anonymization according to feedback from the privacy and utility evaluators. We can see from the results in table 6 that the paradigm of RUPTA can significantly improve the performance of these two baselines. However, due to their being limited to masking or generalizing only entities, their utility preservation performance is still lagged behind RUPTA.
Method Disclosure Risk Utility Preservation
[…]
DB-bio Original 100.00 98.45 99.58 99.68 99.61 99.58 0.0422
DEID-GPT 77.10 79.47 90.82 94.37 92.56 91.22 0.3103
[…]
SD 73.21 73.63 92.27 93.11 92.69 92.96 0.2719
$\text{{SD}}^{*}$ 52.43 54.16 93.98 94.67 94.07 94.10 0.2132
RUPTA (GPT-4) 52.67 53.11 95.58 96.26 95.91 96.02 0.1618
Table 6: Method transferring experiment results on the test set of DB-bio dataset. The top and second performances are highlighted with bold font and underlined, respectively.  denotes the adapted baseline.
Appendix C Customizable Privacy-Utility Tradeoff
In RUPTA, we can manually adjust the privacy-utility tradeoff by setting the maximum of the privacy objective as demonstrated in section 5.2. Besides, as shown in fig. 3, the privacy-utility tradeoff is also changed as the number of optimization steps increases. In this section, we explore the effect of the maximum privacy objective in different optimization steps. We implemented RUPTA using Llama-3-70b here. As shown by the results demonstrated in table 8, the privacy protection level can be actually adjusted in a wider range. Practically, the suitable value of $K$ and $T$ can be empirically chosen by running RUPTA on a validation set and then deploying it in the actual use case. We further conduct experiments to verify the performance of RUPTA is robustness. We list the SR, CS and Accuracy performance of RUPTA on both the validation and test set of DB-bio in table 7. The result is the average of 5 runs on the validation and test set, respectively. We can see that the performance of RUPTA is consistent across the two subsets of DB-bio.
Method SR (s) CS Accuracy
RUPTA-test 64.27 64.11 95.92
RUPTA-val 65.16 64.07 95.46
Table 7: Experiment results of RUPTA on the validation and test set of DB-bio.
Additionally, by comparing the results of SD and RUPTA (K=1) with an optimization step of 1, we observe that RUPTA achieves better utility preservation while maintaining nearly the same level of privacy protection.
Method Disclosure Risk Utility Preservation
@@ Appendix C Customizable Privacy-Utility Tradeoff
Optimization step = 1
RUPTA (K=1) 72.74 73.69 97.12 98.39 97.67 97.11 0.0867
RUPTA (K=5) 68.76 70.23 96.34 97.01 96.15 96.82 0.1121
RUPTA (K=10) 67.32 69.11 96.08 96.56 95.92 96.23 0.1389
Optimization step = 2
RUPTA (K=1) 68.23 69.78 95.12 96.16 95.44 96.09 0.1201
RUPTA (K=5) 65.02 67.93 94.23 95.07 94.89 94.97 0.1608
RUPTA (K=10) 64.67 63.11 94.12 95.23 95.03 94.39 0.1820
Optimization step = 3
RUPTA (K=1) 66.18 68.34 95.39 96.11 95.78 96.06 0.1526
RUPTA (K=5) 65.41 67.14 94.28 95.19 94.88 94.96 0.1599
RUPTA (K=10) 64.23 67.02 94.09 95.10 94.67 94.29 0.1684
@@
Table 8: Privacy-utility tradeoff experiment results in different optimization steps on the test set of the DB-bio dataset. The top and second performances are highlighted with bold font and underlined, respectively.
Appendix D Dataset Settings
Dataset #Train #Validation #Test
DBPedia Classes 1938 243 239
Personal Reddit 318 - 207
Table 9: Statistics of experiment datasets.